# Resource-Efficient Affordance Grounding with Complementary Depth and Semantic Prompts

### v1

@@ I Introduction
The proposed BiT-Align model is verified via a comprehensive variety of experiments and ablation studies on public AGD20K [11] and HICO-IIF [4] datasets. While achieving state-of-the-art affordance grounding performance as shown in Fig. 1, the model’s parameter amount is reduced by $88.8\%$ compared to the current state-of-the-art method [4] with an improvement of $6.0\%$ in the KLD metric, and only $19.2\%$ more than the unimodal baseline model [12], making it highly efficient for practical deployment.
@@ II-A Visual Affordance Grounding
To address the challenges of annotation, many researchers began to draw inspiration from human learning mechanisms and proposed weakly supervised training using interaction videos or photos. Following this, the work in [16] focused on transferring affordance knowledge from third-person to first-person perspectives. With the emergence of self-supervised feature extraction models [17], the study [11] established a training framework for ‘egocentric’ and ‘exocentric’ views and proposed a knowledge transfer mechanism between the two training branches. To further optimize training efficiency and reduce parameters, LOCATE [12] employed clustering to separate foreground and background in views and extract affordance regions, whereas Luo et al. [18] mapped object affordances to human body parts using keypoints. In addition, a piece of recent research in [19] explored the use of compact networks for wearable devices and recent works of [20, 21, 22] conducted experiments in robotic arms, highlights the challenges faced in applying affordance tasks to robotic arms. Current work has established an effective and parameter-efficient affordance training framework, providing highly favorable conditions for further improvements based on this foundation. Based on this, we introduce a variety of modal information to improve performance and control the overall parameter count boost of the network.
@@ II-B Multimodal Affordance Grounding
To enrich training information and enhance the performance of affordance grounding, many studies have begun to integrate textual information into affordance localization. The method in [23] optimized the LOCATE method by using SAM masks and matching them with textual features, while also utilizing LLM chain-of-thought prompting to assist affordance learning. Similarly, the method in [24] used LLM models to obtain textual prompts to aid in affordance learning, and INTRA [25] introduced interaction relationship-guided contrastive learning with LLM, enabling the model to perform without the need for egocentric training branch. Meanwhile, the work in [26] achieved outstanding one-shot open-vocabulary performance by combining CLIP [27] and DINO-V2 [28] encoded features. Additionally, the work in [4] provided a novel perspective by using attention weights to complement textual features. To address the scarcity of multimodal training datasets, some studies, such as AffordanceLLM [29] and GAT [30], leveraged depth estimation [31] to acquire additional modalities for assistance. The introduction of multimodal information improves model accuracy but inevitably increases the number of parameters, raising the costs of model training and inference. Therefore, it is worthwhile to explore methods that introduce additional modalities while maintaining characteristics of a low-parameter framework.
Figure 3: Overview of the proposed BiT-Align framework for image-depth-text affordance mapping. The egocentric branch (ego), exocentric branch (exo), and text branch, form the overall framework of the model. Only the egocentric branch and text branch are retained in the inference stage. The red blocks form the Bypass Prompt Module (BPM), receiving input from depth images and progressively fuses features into the RGB encoder. The designed Text Feature Guidance (TFG) and the Pixel-Text (PT) fusion modules [4] both receive text features and the RGB-D fused features, and then perform fusion in respective methods. In the framework, the weights of the image and text encoders are frozen.
@@ II-C Prompt Learning
The role of prompting in the field of Natural Language Processing (NLP) has been extensively studied and has been proven effective. As NLP methods are increasingly introduced into the vision domain, prompting techniques have gradually gained significant attention from researchers [32]. The research in [33] aimed to leverage prompting to integrate the contextual capabilities of transformers with the prior knowledge of local features in Convolutional Neural Netorks (CNNs). The study presented in [34] employed a typical gourd-shaped adapter for fine-tuning visual transformers, whereas the approach in [35] further combined CNN-like adapter-based prompting methods to generate more compact and robust models for downstream tasks. The work of [36] proposed a parameter-efficient adapter method, and the research in [37] treated multimodal fusion as a form of feature prompting, eliminating the need for multiple encoders, which facilitates faster model inference and training.
@@ III-C Text Feature Guidance
Here, $\mu$ is used as a scaling factor to balance the difference between text and image. In the subsequent steps, we follow the approach outlined in [12]. Building on this, the attention weights obtained from the transformer encoder and self-attention modules can further guide the localization of affordance features. The text features obtained from the text encoder are processed through a simple MLP for a nonlinear mapping, where features are mapped to multiple categories corresponding to the attention heads CLS tokens $\mathcal{H}_{head}^{i}$ ($i{=}0,1,2,...,n$). Subsequently, the MLP outputs a sequence of weights ${h{\in}\mathbb{R}^{c\times n}}$ corresponding to each attention head. The attention heads serve as masks for the output features, and these are multiplied by the original output features to extract attention features $\mathcal{M}^{i}$ from the regions of interest for each attention head. Finally, the resulting attention features are multiplied with the weights output by the MLP, yielding a text-guided attention-weighted feature $\mathcal{F}_{t}$. The obtained features are added to the fused features $\mathcal{F}_{p}$ of the PT module. Among them, the coefficient $\alpha$ is used to balance the weights of the two fusion branches.
@@ III-D Training and Inference
In the testing phase, only the egocentric branch is retained, and RGB, depth, and textual label information are inputted to produce the activation map as the output. The inference branch is further pruned to improve the model’s reasoning speed. Due to physical constraints of robotic manipulators (e.g., end-effector shape and size), practical tool interaction typically occurs at specific sub-regions rather than the entire operable area. Therefore, we represent interaction targets through activation maps instead of segmentation masks. This design concentrates the predicted interaction areas into focused points or peak activation regions with maximum probability density, facilitating precise robotic end-effector positioning for subsequent manipulation tasks.
@@ IV-A Datasets and Data Processing
We mainly use the public Affordance Grounding Dataset (AGD20K) [11] and the HICO-IIF dataset [4] for evaluation. AGD20K includes $20,061$ exocentric and $3,755$ egocentric images. The dataset consists of two subsets: ‘Seen’ and ‘Unseen’ The former contains $36$ affordance labels, while the latter has $25$ affordance labels. The key difference is that the test set in ‘Unseen’ includes categories not present in the training set. This dataset has been widely used to evaluate the performance of affordance methods. HICO-IIF includes $2,885$ exocentric and $1498$ egocentric images. Yet, these datasets do not include depth data. In recent years, the emergence of monocular depth estimation methods with higher accuracy has provided a potential solution to compensate for missing depth data [39, 40, 41]. To assess the role of depth data and the effectiveness of fusion methods, we use Depth Anything V2 [41] to generate corresponding pseudo-depth labels as depth image inputs.
@@ IV-C Qualitative and Quantitative Results
Comparisons with state-of-the-art methods [4, 12, 23, 25] demonstrate that our approach, while inheriting the low-parameter characteristic, integrates depth information and text prompts, achieving the best performance, as shown in Table I. Compared to unimodal methods, our multimodal approach only increases the number of trainable parameters by $19.2\%$, while achieving an $88.7\%$ reduction compared to RGB-D methods. Additionally, we further investigated the advantages of the proposed method in terms of the inference speed. We conducted tests on an NVIDIA RTX3090 GPU. After uniformly resizing the images to $224{\times}224$, we measured the GFLOPs of the model’s inference branch. The inference time was calculated using CUDA calls, starting from the initiation and ending at the completion of the process. With a batch size of $1$, we recorded the inference time over $10$ runs and calculated the average value. The results are shown in Table II. During the inference process, since the BPM module needs to process the input depth data block by block, it inevitably increases the final inference time. However, due to the lightweight structure of the BPM module, the inference time is controlled to approximately 10ms, outperforming publicly available multimodal approaches.
Moreover, we conducted a statistical analysis of the attention heads in DINO based on text input selection, and the results are shown in Fig. 6. It can be observed that the roles of different attention heads indeed exhibit certain distinctions depending on the text labels, which aligns with the findings described in [43]. Since DINO-ViT-S only uses $6$ attention heads and has a smaller scale, the discriminative ability at the parts level is limited. However, the differences observed across two datasets indicate that attention heads still help distinguish foreground from background in different environments.
@@ IV-D Ablation Studies
Finally, due to the simultaneous use of the PT module [4], we conduct additional ablation experiments to evaluate its collaborative performance with the proposed TFG module. The results are shown in Table V. The PT module facilitates text localization but introduces background interference, whereas TFG module counteracts this by dynamically enhancing foreground features through attention-head suppression of background noise, leading to improvements in KLD and SIM.
@@ V Conclusion
In this work, we look into efficient multimodal affordance grounding and put forward the BiT-Align framework using image-depth-text affordance mapping. The framework is constructed with the proposed Bypass Prompt Module (BPM) and Text Feature Guidance (TFG) mechanisms. The proposed BiT-Align effectively integrates depth and semantic information and achieves state-of-the-art performance with only a small number of parameters, paving the way for more resource-efficient and accurate affordance learning and understanding in complex robotic environments. In the future, we intend to further explore interactions between the visual and textual transformer encoders to achieve block-level fusion of textual and visual features.

### v2

@@ I Introduction
The proposed BiT-Align model is verified via a comprehensive variety of experiments and ablation studies on public AGD20K [11] and HICO-IIF [10] datasets. While achieving state-of-the-art affordance grounding performance as shown in Fig. 2, the model’s parameter amount is reduced by $88.8\%$ compared to the current multimodal method [10], with a $6.0\%$ reduction in the KLD metric (lower is better). Compared to the unimodal baseline [12], the proposed model requires only $19.2\%$ additional parameters, making it highly efficient for practical deployment.
@@ II-A Visual Affordance Grounding
To address the challenges of annotation, many researchers began to draw inspiration from human learning mechanisms and proposed weakly supervised training using interaction videos or photos. Following this, the work in [16] focused on transferring affordance knowledge from third-person to first-person perspectives. With the emergence of self-supervised feature extraction models [17], the study [11] established a training framework for ‘egocentric’ and ‘exocentric’ views and proposed a knowledge transfer mechanism between the two training branches. To further optimize training efficiency and reduce parameters, LOCATE [12] employed clustering to separate foreground and background in views and extract affordance regions, whereas Luo et al. [18] mapped object affordances to human body parts using keypoints. In addition, a piece of recent research in [19] explored the use of compact networks for wearable devices, and recent works of [20, 21, 22] conducted experiments in robotic arms, highlighting the challenges faced in applying affordance tasks to robotic arms. Current work has established an effective and parameter-efficient affordance training framework, providing highly favorable conditions for further improvements based on this foundation. On this basis, we introduce deep modal information to improve the understanding of the affordance of the model in the spatial shape.
@@ II-B Multimodal Affordance Grounding
To enrich training information and enhance the performance of affordance grounding, many studies have begun to integrate textual information into affordance localization. The method in [23] optimized the LOCATE method by using SAM masks and matching them with textual features, while also utilizing LLM chain-of-thought prompting to assist affordance learning. Similarly, the method in [24] used LLM models to obtain textual prompts to aid in affordance learning, and INTRA [25] introduced interaction relationship-guided contrastive learning with LLM, enabling the model to perform without the need for an egocentric training branch. Meanwhile, the work in [26] achieved outstanding one-shot open-vocabulary performance by combining CLIP [27] and DINO-V2 [28] encoded features. Additionally, the work in [10] provided a novel perspective by using attention weights to complement textual features. To address the scarcity of multimodal training datasets, some studies, such as AffordanceLLM [29] and GAT [30], leveraged depth estimation [31] to acquire additional modalities for assistance. The introduction of multimodal information improves model accuracy but inevitably increases the number of parameters, raising the costs of model training and inference.
Figure 3: Overview of the proposed BiT-Align framework for image-depth-text affordance mapping. The egocentric branch (ego), exocentric branch (exo), and text branch, form the overall framework of the model. Only the egocentric branch and text branch are retained in the inference stage. The red blocks form the Bypass Prompt Module (BPM), receiving input from depth images and progressively fusing features into the RGB encoder. The designed Text Feature Guidance (TFG) and the Pixel-Text (PT) fusion modules [10] both receive the same text features and RGB-D fused features, and then perform fusion in respective methods. In the framework, the weights of the image and text encoders are frozen.
@@ II-C Prompt Learning
The role of prompting in the field of Natural Language Processing (NLP) has been extensively studied and has been proven effective. As NLP methods are increasingly introduced into the vision domain, prompting techniques have gradually gained significant attention from researchers [32]. The research in [33] aimed to leverage prompting to integrate the contextual capabilities of transformers with the prior knowledge of local features in Convolutional Neural Networks  (CNNs). The study presented in [34] employed a typical gourd-shaped adapter for fine-tuning visual transformers, whereas the approach in [35] further combined CNN-like adapter-based prompting methods to generate more compact and robust models for downstream tasks. The work of [36] proposed a parameter-efficient adapter method, and the research in [37] treated multimodal fusion as a form of feature prompting, eliminating the need for multiple encoders, which facilitates faster model inference and training.
@@ III-C Text Feature Guidance
Here, $\mu$ is used as a scaling factor to balance the difference between text and image. In the subsequent steps, we follow the approach outlined in [12]. Building on this, the attention weights obtained from the transformer encoder and self-attention modules can further guide the localization of affordance features. The text features obtained from the text encoder are processed through a simple MLP for a nonlinear mapping, where features are mapped to multiple categories corresponding to the attention heads CLS tokens $\mathcal{H}_{head}^{i}$ ($i{=}0,1,2,...,n$). Subsequently, the MLP outputs a sequence of weights ${h{\in}\mathbb{R}^{c\times n}}$ corresponding to each attention head. The attention heads serve as masks for the output features, and these are multiplied by the original output features to extract attention features $\mathcal{M}^{i}$ from the regions of interest for each attention head. Finally, the resulting attention features are multiplied by the weights output by the MLP, yielding a text-guided attention-weighted feature $\mathcal{F}_{t}$. The obtained features are added to the fused features $\mathcal{F}_{p}$ of the PT module. Among them, the coefficient $\alpha$ is used to balance the weights of the two fusion branches.
@@ III-D Training and Inference
In the testing phase, only the egocentric branch is retained, and RGB, depth, and textual label information are inputted to produce the activation map as the output. Since the exocentric branch encodes explicit knowledge, the model should learn to reason about possible actions solely from the egocentric branch, which lacks human actions. Due to robotic manipulator constraints (e.g., end-effector shape/size), tool interaction typically occurs at specific sub-regions rather than the entire operable area. Thus, we represent interaction targets via activation maps (not segmentation masks), concentrating predictions into focused peak regions with maximum probability density to guide precise end-effector positioning.
@@ IV-A Datasets and Data Processing
We mainly use the public Affordance Grounding Dataset (AGD20K) [11] and the HICO-IIF dataset [10] for evaluation. AGD20K includes $20,061$ exocentric and $3,755$ egocentric images. The dataset consists of two subsets: ‘Seen’ and ‘Unseen’. The former contains $36$ affordance labels, while the latter has $25$ affordance labels. The key difference is that the test set in ‘Unseen’ includes categories not present in the training set. This dataset has been widely used to evaluate the performance of affordance methods. HICO-IIF includes $2,885$ exocentric and $1498$ egocentric images. Yet, these datasets do not include depth data. In recent years, the emergence of monocular depth estimation methods with higher accuracy has provided a potential solution to compensate for missing depth data [39, 40, 41]. To assess the role of depth data and the effectiveness of fusion methods, we use Depth Anything V2 [41] to generate corresponding pseudo-depth labels as depth image inputs.
@@ IV-C Qualitative and Quantitative Results
Comparisons with state-of-the-art methods [10, 12, 23, 25] demonstrate that our approach, while inheriting the low-parameter characteristic, integrates depth information and text prompts, achieving the best performance, as shown in Table I. Compared to unimodal methods, our multimodal approach only increases the number of trainable parameters by $19.2\%$, while achieving an $88.7\%$ reduction compared to RGB-D methods. Additionally, we further investigated the advantages of the proposed method in terms of the inference speed. We conducted tests on an NVIDIA RTX3090 GPU. After uniformly resizing the images to $224{\times}224$, we measured the GFLOPs of the model’s inference branch. The inference time was calculated using CUDA calls, starting from the initiation and ending at the completion of the process. With a batch size of $1$, we recorded the inference time over $10$ runs and calculated the average value. The results are shown in Table II. While the BPM module’s block-wise processing of depth data increases inference time, its lightweight design limits this overhead to just $10.0$ ms faster than the public multimodal approach [10] which takes $12.8$ ms.
Moreover, we conducted a statistical analysis of the attention heads in DINO based on text input selection, and the results are shown in Fig. 6. It can be observed that the roles of different attention heads indeed exhibit certain distinctions depending on the text labels, which aligns with the findings described in [43]. Since DINO-ViT-S only uses $6$ attention heads and has a smaller scale, the discriminative ability at the parts level is limited. However, the differences observed across the two datasets indicate that attention heads still help distinguish foreground from background in different environments.
@@ IV-D Ablation Studies
Finally, due to the simultaneous use of the PT module [10], we conduct additional ablation experiments to evaluate its collaborative performance with the proposed TFG module. The results are shown in Table V. The PT module facilitates text localization but introduces background interference, whereas the TFG module counteracts this by dynamically enhancing foreground features through attention-head suppression of background noise, leading to improvements in KLD and SIM.
@@ V Conclusion
In this work, we look into efficient multimodal affordance grounding and put forward the BiT-Align framework using image-depth-text affordance mapping. The framework is constructed with the proposed Bypass Prompt Module (BPM) and Text Feature Guidance (TFG) mechanisms. The proposed BiT-Align effectively integrates depth and semantic information and achieves state-of-the-art performance with only a small number of parameters, paving the way for more resource-efficient and accurate affordance learning and understanding in complex robotic environments.
In the future, we intend to further evaluate the framework’s robustness to depth noise and domain shifts, while extending its application to more challenging real-world scenarios. Additionally, we will investigate deeper interactions between visual and textual encoders to enhance affordance reasoning.